The following transactions occurred during 2017, Assume that depreciation of 10% per year is charged on all machinery and 5% per year on buildings, on a straight-line basis, with no estimated salvage value. Depreciation is charged for a full year on all fixed assets acquired during the year, and no depreciation is charged on fixed assets disposed of during the year. Jan. 30 A building that cost $187,440 in 2000 is torn down to make room for a new building. The wrecking contractor was paid Mar. 10 Machinery that was purchased in 2010 for $22,720 is sold for $4,118 cash, f.о.b. purchasers plant. Freight of S426 is Mar. 20 A gear breaks on a machine that cost $12,780 in 2009. The gear is replaced at a cost of $2,840. The replacement does May 18 A special base installed for a machine in 2011 when the machine was purchased has to be replaced at a cost of $7,242 and was permitted to keep all materials salvaged. paid on the sale of this machinery. not extend the useful life of the machine but does make the machine more efficient. $7,810 because of defective workmanship on the original base. The cost of the machinery was $20,164 in 2011. The cost of the base was $4,970, and this amount was charged to the Machinery account in 2011. One of the buildings is repainted at a cost of $9,798. It had not been painted since it was constructed in 2013. June 23 Prepare general journal entries for the transactions. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter O for the amounts.) Date Account Titles and Explanation Debit Credit 1/303/10 5/185/18
Answer #1

SOLUTION

Journal entries-

Date Accounts titles and Explanation Debit ($) Credit ($)
Jan.10 Accumulated depreciation 159,324
Loss on disposal (Balancing figure) 35,358
Building 187,440
Cash 7,242
Mar.10 Cash 3,692
Accumulated depreciation 15,904
Loss on sale (Balancing figure) 3,124
Machinery 22,720
Mar.20 Machinery 2,840
Cash 2,840
May 18 Machinery 7,810
Accumulated depreciation 12,096
Loss on disposal (Balancing figure) 8,068
Machinery 20,164
Cash 7,810
June 23 Repairs and maintenance expense 9,798
Cash 9,798

1. Jan.30

Depreciation for one year = $187,440*5% = 9,372

Accumulated depreciation = $9,372 * 17 years = $159,324

2. Mar.10

Depreciation for one year = $22,720*10% = 2,272

Accumulated depreciation = $2,272*7 years = $15,904

Cash = $4,118 - $426 = $3,692

3. May 18

Depreciation for one year = $20,164*10% = 2,016

Accumulated depreciation = $2,016*6 years = $12,096